# An efficient discrete unified gas kinetic scheme for strongly inhomogeneous fluids at the nanoscale

### v1

@@ Abstract
The kinetic model with multiple integral terms based on the Enskog-Vlasov(EV) equation is widely employed to describe the inhomogeneous fluids at the nanoscale. However, previous studies have mainly focused on one-dimensional cases, partly due to the significant computational cost $O(NN_{\sigma})$ associated with direct computation of integrals, where $N$ is the number of cells in the flow field and $N_{\sigma}$ is the number of cells in a cube with a side length equal to the molecular diameter $\sigma$. In this study, we propose a discrete unified gas kinetic scheme (DUGKS) with efficient numerical strategies for integrals to overcome the inefficiency of the direct method, reducing the computational cost to $O(N)$. Both accuracy and efficiency of the proposed DUGKS are assessed through several test cases, including static fluid structures and force-driven flow dynamics in parallel plate channels. As example applications, pressure-driven flow between two flat plates and force-driven flow in a square duct are investigated to highlight distinctive phenomena at the nanoscale.
@@ 1 Introduction
In this study, we will develop an improved DUGKS by evaluating the spatial integrals with much reduced grid points without sacrificing the accuracy, thus the computational efficiency is significantly improved. The remainder of this paper is organized as follows. Section  2 provides a review of the strongly inhomogeneous kinetic model at the nanoscale. Next, the original DUGKS for the kinetic model is reviewed in Section  3. In Section  4, we present the detailed formulations of the proposed DUGKS, including the treatments of the integrals. In Section  5, we assess the performance of the proposed DUGKS through several numerical tests. Section  6 investigates two-dimensional flows. Finally, concluding remarks are presented in Section  7.
@@ 5.2 Force-driven flow between two flat plates
5.3 Comparison of computational complexity
The computational costs of the proposed DUGKS and the original DUGKS with direct computations for integrals are studied in force-driven flows. The spatial domain is discretized with $N_{\sigma x}=\{20,40\}$ and $N_{\sigma y}=\{50,100,150,200\}$, which are necessary to ensure the grid independence of the results due to variations in fluid properties along both the $x$- and $y$-direction. The speedup for various values of $N_{\sigma}=N_{\sigma x}N_{\sigma y}$ is presented in Fig. 7. As expected, the speedup $Sp$ is proportional to the $N_{\sigma}$. For a fixed $N_{\sigma y}$, the speedup at $N_{\sigma x}=40$ is twice that at $N_{\sigma x}=20$. When $N_{\sigma x}=40$, the speedup can reach up to two orders of magnitude for $N_{\sigma y}>100$, whereas for $N_{\sigma x}=20$, $N_{\sigma y}$ needs to be increased to $200$ to achieve the same level of speedup. As $N_{\sigma z}$ increases in three-dimensional flows, speedup $Sp$ is expected to further increase to three orders of magnitude.
@@ 6.1 Pressure-driven flow between two flat plates
We now consider pressure-driven flow in a channel of width $H=10.26\sigma$ and length $L=5H$. The channel connects two reservoirs maintained at different pressures. The average number density in the outlet reservoir is fixed at $n_{out}=0.10\,\sigma^{-3}$, while the inlet reservoir value $n_{in}$ is varied to control the pressure ratio and investigate its effect on the flow. The physical domain is discretized on a uniform Cartesian mesh with $N_{\sigma y}=200$ and $N_{\sigma x}=20$. In velocity space, we employ an $8\times 8$ half-range Gauss-Hermite quadrature.
@@
Based on numerical experiments, we impose nanoscale pressure boundary conditions using a non-equilibrium extrapolation scheme [wu2016discrete, zhao2002non] to ensure a smooth transition of macroscopic fields from the interior to the boundary.
@@
Fig. 8 shows the density and streamwise velocity profiles for outlet number densities $n_{out}=0.15\,\sigma^{-3}$ and $0.40\,\sigma^{-3}$. The density and streamwise velocity are approximately linear along the flow direction at $n_{out}=0.15\,\sigma^{-3}$, whereas they become strongly nonlinear at $n_{out}=0.40\,\sigma^{-3}$. Fig. 9 compares the pressure-driven flow with the corresponding force-driven case, with the external force evaluated from Eq. (46). While the density agrees well, the velocity shows noticeable discrepancies owing to a non-negligible deviation from a linear streamwise pressure distribution at the nanoscale. This conflicts with the assumption in Eq. (46) of a uniform streamwise pressure gradient, i.e., a linear pressure profile. Under isothermal conditions (so that $p\propto n$), we quantify the pressure deviation as
$\delta p(x)=\frac{p(x)-p_{l}(x)}{p_{out}}=\frac{n(x)-n_{l}(x)}{n_{out}},\quad n_{l}(x)=n_{in}+\frac{x}{L}(n_{out}-n_{in}\bigr),$ (49)
where $n(x)$ denotes the cross-sectional average number density. Fig. 10 shows that the pressure deviation $\delta p$ increases with the pressure ratio $p_{in}/p_{out}$. We further observe that, at the nanoscale, the streamwise-averaged velocity ratio $R_{u}$ between pressure-driven and force-driven cases is approximately proportional to the maximum pressure deviation $\delta p_{\max}=\max_{x}\delta p(x)$ over a range of pressure ratios, as shown in Fig. 11. These results demonstrate that pressure-driven and force-driven flows are not generally equivalent at the nanoscale.
Figure 10: Pressure deviation $\delta p(x)$ along the flow direction for $p_{in}/p_{out}=1.5,\,2.0,\,2.5,\,3.0,\,3.5,$ and $4.0$. From bottom to top, the curves follow the same order.
Figure 11: Streamwise-averaged velocity ratio $R_{u}$ versus maximum pressure deviation $\delta p_{\max}$ across pressure ratios $p_{in}/p_{out}\in\{1.5,\,2.0,\,2.5,\,3.0,\,3.5,\,4.0\}$.
@@ 6.2 Force-driven flow in a square duct
Consider force-driven flow in square ducts with widths $H=10.26\sigma$ and $H=5.13\sigma$. In both cases, the cross-sectional average number density is fixed at $n_{0}=0.30\,\sigma^{-3}$. A constant external force $\bm{F}_{i}=(0,0,F_{z})$ with $F_{z}=0.032\,\epsilon_{ff}/\sigma$ is applied to each molecule along the $z$-direction. The physical domain is discretized on a uniform Cartesian mesh with $N_{\sigma x}=N_{\sigma y}=200$ cells. In velocity space, we employ an $8\times 8\times 8$ half-range Gauss-Hermite quadrature.
@@
with $\nu$ the kinematic viscosity. For $H=10.26\sigma$, the Knudsen minimum is observed at $Kn\approx 0.1$; for $H=5.13\sigma$, the Knudsen minimum disappears and the mass flow rate $Q_{n}$ increases monotonically with $Kn$.
@@ 7 Conclusions
Although the proposed kinetic model can accurately describe inhomogeneous fluids at the nanoscale, the direct computations of multiple integral terms associated with nonlocal gradient, average density, and mean-field force lead to a significant computational cost $O(NN_{\sigma})$ in the original DUGKS. In this study, we propose a novel DUGKS that incorporates efficient numerical schemes for these integrals, reducing the computational cost to $O(N)$. A new difference method is introduced to replace the integral method for nonlocal gradients, and an adaptive volume-averaging method for average density is proposed to accurately describe two-dimensional flows. The average density and the mean-field potential are computed on a coarse scale, where fewer points are used to compute the mean-field potential as distance increases due to the rapid decay of attractive potential in space.
The accuracy of the proposed DUGKS has been validated through numerical tests in static fluid structures and flow dynamics. The speedup relative to the original DUGKS has been rigorously calculated. It is shown that the speedup is proportional to $N_{\sigma}$ and can reach up to two orders of magnitude in two-dimensional flows. These results demonstrate that the proposed DUGKS is a powerful numerical tool for nanoscale kinetic models. We further investigate two-dimensional flows to reveal the exotic phenomena at the nanoscale. In pressure-driven channel flow, we find that pressure-driven and force-driven flows are not generally equivalent because the streamwise pressure gradient is non-uniform. In force-driven square duct flow, density peaks are observed near the corners due to the combined attraction from adjacent walls.
@@ Acknowledgments
The computation is performed in the HPC Platform of Huazhong University of Science and Technology.

### v2

@@ Abstract
We develop an efficient discrete unified gas kinetic scheme (DUGKS) to solve a kinetic model for strongly inhomogeneous fluid flows at the nanoscale. The proposed DUGKS adopts efficient numerical strategies to evaluate the multiple nonlocal integrals involved in the kinetic equation, reducing the computational cost from $O(NN_{\sigma})$ for the direct evaluation to $O(N)$, where $N$ and $N_{\sigma}$ are the number of cells in the flow domain and integral domain, respectively. The accuracy and efficiency of the proposed DUGKS are assessed through several test cases, including static fluid structures and force-driven flow in nano slits. The present results agree well with those from Monte Carlo and molecular dynamics simulations, as well as with those from the original DUGKS with direct evaluation of the integrals. Moreover, the speedup relative to the original DUGKS can reach up to two orders of magnitude. As example applications, we investigate two-dimensional pressure-driven flow between parallel plates and force-driven flow in a square duct, revealing distinctive nanoscale features such as the non-equivalence of pressure- and force-driven flows and the formation of density peaks in the square duct corners.
@@ 1 Introduction
In this study, we develop an improved DUGKS by evaluating the spatial integrals with much reduced grid points without sacrificing the accuracy, thus the computational efficiency is significantly improved. The remainder of this paper is organized as follows. Section  2 provides a review of the strongly inhomogeneous kinetic model at the nanoscale. Next, the original DUGKS for the kinetic model is reviewed in Section  3. In Section  4, we present the detailed formulations of the proposed DUGKS, including the treatments of the integrals. In Section  5, we assess the performance of the proposed DUGKS through several numerical tests. Section  6 investigates two-dimensional flows. Finally, concluding remarks are presented in Section  7.
@@ 5.2 Force-driven flow between two flat plates
5.3 Comparison of computational efficiency
The computational costs of the proposed DUGKS and the original one are compared for force-driven flows. The spatial domain is discretized with $N_{\sigma x}=\{20,40\}$ and $N_{\sigma y}=\{50,100,150,200\}$, which are necessary to ensure the grid independence of the results due to variations in fluid properties along both the $x$- and $y$-directions. The speedup is defined as
$Sp=\frac{T_{o}}{T_{p}},$ (47)
where $T_{o}$ and $T_{p}$ denote the runtimes for $1000$ iteration steps using the original DUGKS and the present DUGKS, respectively. The corresponding values of $Sp$ are shown in Fig. 7 as a function of $N_{\sigma}=N_{\sigma x}N_{\sigma y}N_{\sigma z}$, with $N_{\sigma z}=1$. As expected, the speedup is proportional to the $N_{\sigma}$. For a fixed $N_{\sigma y}$, the speedup at $N_{\sigma x}=40$ is twice that at $N_{\sigma x}=20$. When $N_{\sigma x}=40$, the speedup can reach up to two orders of magnitude for $N_{\sigma y}>100$, whereas for $N_{\sigma x}=20$, $N_{\sigma y}$ needs to be increased to $200$ to achieve the same level of speedup. For three-dimensional flows with $N_{\sigma z}>1$, the speedup is expected to increase further to three orders of magnitude.
@@ 6.1 Pressure-driven flow between two flat plates
We first consider pressure-driven flow in a channel of width $H=10.26\sigma$ and length $L=5H$. The channel connects two reservoirs maintained at different pressures. The average number density in the outlet reservoir is fixed at $n_{out}=0.10\,\sigma^{-3}$, while the number density in the inlet reservoir $n_{in}$ is varied to control the number density ratio $n_{in}/n_{out}$ and investigate its influence on the flow. The physical domain is discretized on a uniform Cartesian mesh with $N_{\sigma y}=200$ and $N_{\sigma x}=20$. The velocity space is discretized by the abscissa of an $8\times 8$ half-range Gauss-Hermite quadrature.
@@
Then, we impose the nanoscale pressure boundary conditions using a non-equilibrium extrapolation scheme [wu2016discrete, zhao2002non] to ensure a smooth transition of macroscopic fields from the interior to the boundaries.
@@
Fig. 8 shows the density and streamwise velocity profiles for the cases of $n_{out}=0.15\,\sigma^{-3}$ and $0.40\,\sigma^{-3}$. The density and streamwise velocity are approximately linear along the flow direction at $n_{out}=0.15\,\sigma^{-3}$, whereas they become strongly nonlinear at $n_{out}=0.40\,\sigma^{-3}$. Fig. 9 compares the pressure-driven flow with the corresponding force-driven case, with the external force evaluated from Eq. (46). While the density agrees well, the velocity shows noticeable discrepancies owing to a non-negligible deviation from a linear streamwise density distribution at the nanoscale. This conflicts with the assumption in Eq. (46) of a uniform streamwise density gradient, i.e., a linear density profile. To quantify the deviation from this linear profile, we define
$\delta n(x)=\frac{n(x)-n_{l}(x)}{n_{out}},\quad n_{l}(x)=n_{in}+\frac{x}{L}(n_{out}-n_{in}\bigr),$ (50)
where $n(x)$ denotes the cross-sectional average number density and $n_{l}(x)$ is the reference linear profile connecting $n_{in}$ at $x=0$ and $n_{out}$ at $x=L$. Fig. 10 shows that the density deviation $\delta n$ increases with the density ratio $n_{in}/n_{out}$. We further observe that, at the nanoscale, the streamwise-averaged velocity ratio $R_{u}$ between pressure-driven and force-driven cases varies approximately linearly with the maximum number density deviation $\delta n_{\max}=\max_{x}\delta n(x)$ over a range of density ratios, as shown in Fig. 11. These results demonstrate that pressure-driven and force-driven flows are not generally equivalent at the nanoscale [guan2022evaluation].
Figure 10: Density deviation $\delta n(x)$ along the flow direction for $n_{in}/n_{out}=1.5,\,2.0,\,2.5,\,3.0,\,3.5,$ and $4.0$. From bottom to top, the curves follow the same order.
Figure 11: Streamwise-averaged velocity ratio $R_{u}$ versus maximum density deviation $\delta n_{\max}$ across density ratios $n_{in}/n_{out}\in\{1.5,\,2.0,\,2.5,\,3.0,\,3.5,\,4.0\}$.
@@ 6.2 Force-driven flow in a square duct
We now consider fully developed force-driven flow in infinitely long straight square ducts with widths $H=10.26\sigma$ and $H=5.13\sigma$. The duct axis is aligned with the $z$-direction and the cross-section lies in the $x$-$y$ plane. Owing to the fully developed assumption, all macroscopic quantities are uniform along the $z$-direction, so the problem is effectively two-dimensional in physical space and the simulation is performed on the $x$-$y$ cross-section. In both cases, the cross-sectional average number density is fixed at $n_{0}=0.30\,\sigma^{-3}$. A constant external force $\bm{F}_{i}=(0,0,F_{z})$ with $F_{z}=0.032\,\epsilon_{ff}/\sigma$ is applied to each molecule along the $z$-direction to drive the flow. The physical domain is discretized by a uniform Cartesian mesh with $N_{\sigma x}=N_{\sigma y}=200$ cells. Although the spatial domain is treated as two-dimensional, the velocity space remains three-dimensional and is discretized using an $8\times 8\times 8$ half-range Gauss–Hermite quadrature.
@@
with $\nu$ the kinematic viscosity. For $H=10.26\sigma$, the Knudsen minimum is observed at $Kn\approx 0.1$; for $H=5.13\sigma$, the Knudsen minimum disappears and the mass flow rate $Q_{n}$ increases monotonically with $Kn$. It is interesting that the disappearance of the Knudsen minimum has also been observed in previous studies [guo2006generalized].
@@ 7 Conclusions
The DUGKS developed in Ref. [shan2020discrete] is capable of capturing the flow behaviors of nano-confined strongly inhomogeneous fluids. However, direct computation of multiple integral terms associated with nonlocal gradient, average density, and mean-field force leads to a significant computational cost $O(NN_{\sigma})$, where $N$ and $N_{\sigma}$ represent the number of cells in the entire physical space and integral domain. As a result, previous studies have been limited to one-dimensional problems. In this study, we propose a novel DUGKS that incorporates efficient numerical schemes for these integrals, reducing the computational cost to $O(N)$. A new difference method is introduced to replace the integral method for nonlocal gradients, and an adaptive volume-averaging method for average density is proposed to accurately describe two-dimensional flows. The average density and the mean-field potential are computed on a coarse scale, where fewer points are used to compute the mean-field potential as distance increases due to the rapid decay of attractive potential in space.
The accuracy of the proposed DUGKS has been validated through numerical tests in static fluid structures and flow dynamics. The speedup relative to the original DUGKS has been rigorously calculated. It is shown that the speedup is proportional to $N_{\sigma}$ and can reach up to two orders of magnitude in two-dimensional flows. These results demonstrate that the proposed DUGKS is a powerful numerical tool for simulating nanoscale flows.
We further investigate two-dimensional flows to reveal the exotic phenomena at the nanoscale. In the pressure-driven channel flow, we find that pressure-driven and force-driven flows are generally not equivalent because the density distribution along the streamwise direction is nonlinear. In the force-driven square duct flow, density peaks are observed near the corners due to the combined attraction from adjacent walls. In future work, we plan to apply the proposed DUGKS to complex three-dimensional and thermal flows.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 12472290). The computation is performed in the HPC Platform of Huazhong University of Science and Technology.